QUESTION 4

Nonuniform Inputs, Weighted Average

Carter Inc. had the following production and cost information for its fabrication department during April (with materials added at the beginning of the fabrication process):

Production:
Units in process, April 1, 50% complete with respect to conversion 10,000
Units completed 65,200
Units in process, April 30, 60% complete 12,000
Costs:
Work in process, April 1:
Materials $40,980
Conversion costs 14,600
   Total $55,580
Current costs:
Materials $125,000
Conversion costs 210,000
   Total $335,000

Carter uses the weighted average method.

Required:

1. Prepare an equivalent units schedule.

Carter Inc.
Schedule of Equivalent Units
(Weighted Average Method)
Materials Conversion
Units completed
Add: Units in ending WIP x Percent complete
× %
× %
Equivalent units of output

2. Calculate the unit cost. Round all intermediate calculations and final answer to the nearest cent.
$per unit

3. Calculate the cost of units transferred out and the cost of EWIP. In your calculations round unit cost to the nearest cent and your final answers to the nearest dollar.

Cost transferred out $
Cost of ending work in process
Answer #1

1. Prepare an equivalent units schedule.

Carter Inc.
Schedule of Equivalent Units
(Weighted Average Method)
Materials Conversion
Units completed 65200 65200
Add: Units in ending WIP x Percent complete
12000 × 100% 12000
12000 × 60% 7200
Equivalent units of output 77200 72400

2. Calculate the unit cost. Round all intermediate calculations and final answer to the nearest cent.

Material cost per unit = (40980+125000)/77200 = 2.15

Conversion Cost per unit = (14600+210000)/72400 = 3.10

$5.25 per unit

3. Calculate the cost of units transferred out and the cost of EWIP. In your calculations round unit cost to the nearest cent and your final answers to the nearest dollar.

Cost transferred out (65200*5.25) $342300
Cost of ending work in process (12000*2.15+7200*3.1) 48120